# Language Models and Logic Programs for Trustworthy Financial Reasoning

### v2

@@ Abstract
According to the United States Internal Revenue Service, “the average American spends \$270\\$270 and 13 hours filing their taxes”. Even beyond the U.S., tax filing requires complex reasoning, combining application of overlapping rules with numerical calculations. Because errors can incur costly penalties, any automated system must deliver high accuracy and auditability, making modern large language models (LLMs) poorly suited for this task. We propose an approach that integrates LLMs with a symbolic solver to calculate tax obligations. We evaluate variants of this system on the challenging StAtutory Reasoning Assessment (SARA) dataset, and include a novel method for estimating the cost of deploying such a system based on real-world penalties for tax errors. We further show how combining up-front translation of plain-text rules into formal logic programs, combined with intelligently retrieved exemplars for formal case representations, can dramatically improve performance on this task and reduce costs to well below real-world averages. Our results demonstrate the promise and economic feasibility of neuro-symbolic architectures for increasing equitable access to reliable tax assistance.
@@ Introduction
Accuracy in tax filing is essential. Over-reported income or missed deduction opportunities lead to unnecessary over-payment, while under-reporting may result in penalties, interest, and potential legal consequences. In the United States, the costs of inaccuracies affect lower income communities more significantly, in part because these groups offer the Internal Revenue Service (IRS) high audit success rates (Black et al. 2022; Elzayn et al. 2025). However, these audits deliver a modest return on investment compared to audits of wealthier taxpayers, so there is an opportunity to better align community and institutional interests through improved tax advice to lower income taxpayers (Boning et al. 2024).
@@
Our experiments demonstrate two key findings. First, whereas frontier reasoning models outperform non-reasoning models at both directly solving and at parsing case and statute text into the symbolic solver, non-reasoning models consistently outperform their reasoning counterparts when given gold symbolic representations of statutes and of their application to similar cases. Second, we show that by adding additional refusal criteria through a symbolic solver and self-checking, the expected costs of deploying such a system in the real world could be brought down to less than 20% of the average cost for an American to file their taxes. Our results indicate the promise of neuro-symbolic architectures for expanding access to trustworthy and reliable tax expertise.
@@
Several programming languages have been designed to represent and facilitate logical reasoning. Prolog is a declarative programming language for representing and reasoning over knowledge, with roots in first order logic. A programmer defines rules using Horn clauses (Horn 1951) and facts by declaring which rules apply to entities, thus populating a knowledge base. Subsequently, this knowledge base can be queried by defining a ‘goal’, which launches computation in the form of a backward-chaining search attempting to prove that the goal holds a certain value (Wielemaker et al. 2010). Prolog has been used since the early days of legal AI, where it has formed the backbone of legal expert systems because its declarative syntax keeps knowledge base entries for rules human-readable while powerfully representing the reasoning around which legal questions revolve (Sherman 1989). Efforts in countries like the United Kingdom (Sergot et al. 1986), Canada (Sherman 1987), and the United States (Kant et al. 2025) have leveraged this capacity to encode legal rules in executable formal logic. Related languages have also been used in the legal domain, such as Answer-set Programming (Gelfond and Lifschitz 1988; Morris 2020), Datalog (Ceri et al. 1989; Huang et al. 2021), and Catala (Merigoux et al. 2021a; Merigoux 2023). More broadly, hierarchical templates are a popular tool for evaluation of legal reasoning (Hou et al. 2024). We focus on the SARA dataset (Holzenberger et al. 2020), which encodes statutes and cases into Prolog logic programs to show how a symbolic expert system can perfectly solve a task which large language models struggle to complete.
@@
We evaluate our methods on the SARA dataset, which tests the ability of language models to do statutory reasoning about the United States Tax Code (Holzenberger et al. 2020). This dataset is included in the popular aggregate benchmark LegalBench (Guha et al. 2023), and was used in the GPT-4 product launch to highlight the model’s superior reasoning capacity (Blair-Stanek et al. 2024). The SARA dataset consists of 9 sections from the US federal tax code which have been moderately edited to make them self-contained and unambiguous. These manipulations allow the dataset to serve as a self-contained and solvable task for language models that lack live internet access and human-like abilities to process ambiguity (Jurayj et al. 2022; Stengel-Eskin et al. 2024). These statute sections are accompanied by 376 hand-crafted cases to test understanding of these statutes, each containing a question about a person’s tax obligation. Each statute and case has been manually translated into Prolog, which allows them to be trivially solved using the language’s powerful execution engine to resolve queries about cases. This Prolog is defined using neo-Davidsonian event semantics (Davidson 1966), thus categorizing each event as one of 61 possible predicates onto which various arguments are attached. Of the 376 cases and corresponding questions, 276 require binary judgments about whether a statute applies to a given case, and 100 require numerical judgments about how much tax a person owes in a given year. We focus on these 100 tax cases because of their increased difficulty, and because a trivial baseline of always guessing a single answer delivers poor performance at predicting the numerical output.
Zero-Shot Solving from Statutory Text
@@
Calculation by Parsing for a Symbolic Solver
To extend the direct solution approach, we augment the language model with a symbolic solver. Here, a model is given the plain text of the statutes as in the direct calculation case. It is instructed to generate a Prolog program which encodes the relevant rules and facts necessary to compute the person in question’s tax obligation. The symbolic solver ingests a set of rules and facts in Prolog, and is invoked to execute a query. The execution of this Prolog program offers a straightforward mechanism for refusal: if the program fails to execute into the proper format or hangs beyond a pre-allocated time limit (10 seconds), the system is considered to have refused to answer.
@@
Table 2: Results of different methods without gold statutes. Models in the same family have the same base model (or seem most likely to, in the case of closed-weights models). Note that “break-even price” measures only the costs of failures and abstentions, and does not include inference costs. For each model, the approach that delivered the lowest break-even price is shown in bold. The top two rows show the break-even price of trivial systems, which always defer to an expert or which always tell a person not to pay any taxes. Errors represent a 90% confidence interval. The lowest break-even price method for each model family is in bold.
@@
We run experiments across four model families of different sizes, three of which are open-weight models. The bases for these models are: Qwen2.5 32B (Qwen et al. 2025), Llama 3.3 70B (Grattafiori et al. 2024), DeepSeek-V3 671B (DeepSeek-AI et al. 2025b), and OpenAI’s GPT-4.1 (OpenAI 2025a, b), each of which has an instruction-tuned version designed for common chat applications, and a reasoning version optimized to expend additional inference-time compute to solve harder problems. The full list of models is included in table 1. We run auxiliary experiments using GPT-5, but we do not conduct the same chat vs. reasoning comparison because this product appears to be an integrated system containing several models, and therefore is not as analogous to these other comparisons (Zhang et al. 2025). The reasoning model for three of these pairings stem from the DeepSeek R1 project (DeepSeek-AI et al. 2025a), where strong base models were fine-tuned to generate long chains-of-thought that help them solve harder quantitative reasoning problems. Although there is no formal documentation stating that OpenAI’s GPT-4.1 and o3 models are derived from the same base model, the proximity of the two model’s launch dates and their identical per-token pricing schemes suggest that these two models have a similar relationship to the other model pairs we explore.
@@
We further ask how effectively these methods can serve to improve each other’s selectivity, by using comparisons between different solution methods to expend additional compute to help determine whether an answer should be trusted (Wang et al. 2023; Stengel-Eskin and Van Durme 2023; Jurayj et al. 2025). In these settings, an answer is only accepted if it is reached via two independent reasoning processes: two chains-of-thought and answers (either directly calculated obligations or Prolog programs) are sampled from the same model. When self-checking using the same method (for instance, “Parsed + Parsed”), these answers are conditioned on the same prompt and context. In the parsing-based approaches, this can be considered a more stringent version of the existing refusal system which rejects questions for which the parsed solution does not execute, by additionally deferring to a tax professional where a combination of attempts do not reach a consensus answer. We test the effectiveness of each combination of reasoning processes for each model to show where additional selectivity can further improve performance.
@@
We posit that the deployment of an automated tax advisor system should require some accountability for the organization deploying it. To provide a realistic estimate of the costs of employing the methods we outline above draw from the US Internal Revenue Code (IRC) $\S 6662$, which imposes a penalty of $20\%$ of the amount underpaid for “substantial understatement of income tax”. The cases in SARA are all personal or household tax cases, so a tax filing is considered a “substantial understatement” if they are less than the maximum of $10\%$ of the actual amount owed or \$5,000\\$5,000. We finally impose a penalty of \$270\\$270 for refusing to answer, to model the cost of the personal or professional time required to complete one’s taxes (Internal Revenue Service 2025).
@@
$\text{cost}=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}-(\Delta y_{i}),&\Delta y_{i}<0,\\[6.0pt] 0.2\,(\Delta y_{i}),&\Delta y_{i}>\max\!\bigl{(}\$5,000,\;0.1\cdot y_{i}\bigr{)},\\[6.0pt] $\$270$,&\text{if refused},\\[6.0pt] $\$0$,&\text{otherwise}.\par\end{cases}$
@@
Exemplars Successes Failures Break-Even Price
Ranked 87 8 \$247.99\\$247.99
Random 71 24 \$5459.25\\$5459.25
Table 3: Intelligently re-ranking few-shot exemplars improves performance. We measure the advantage of intelligent versus random exemplar re-ranking on GPT-4.1.
Few-Shot Parsing using Gold Statutes
We additionally consider what advantage could be given by offering language model parsers access to gold symbolic representations of the statutes, alongside demonstrations of cases applying the rules from these statutes, building on previous work showing how demonstrations can aid in semantic parsing (Shin and Van Durme 2022; Spiegel et al. 2024). Notably, the gold parsed cases all reference the same manually-translated Prolog representation of the statutes. As such, these manual translations formalize certain types of facts in one particular way where several may be viable, given alternative representations of the rules which are plausible but not implemented in practice. Thus, when using already parsed cases as few-shot examples, the execution engine must have access to this particular formalization of the rules as well.
To identify the most salient exemplar cases, we apply information retrieval systems which can follow instructions (Weller et al. 2025a), directing a retrieval system to rank cases based on how similar the logical structure of the case’s text is to the case at hand. For each case, we instruct a lightweight reasoning model (OpenAI o4-mini) to rank the other 99 cases, following recent work showing the effectiveness of test-time scaling for reranking documents (Weller et al. 2025b; Yang et al. 2025). As few-shot examples, we provide the 5 most relevant cases and their gold Prolog translations in-context for the language model to condition its parse on. These cases can be analogized to precedents, because our tax calculation agent uses them to understand how terms from the statutes are applied in practice. We note that this use of the term ‘precedent’ is informal and does not refer to the common-law practice of binding precedent. Rather, these retrieved precedent cases are more analogous to ‘persuasive’ precedent, which might help a court understand how terms have been applied in the past but is not itself new law which future courts must apply (Kozel 2014).
To test the value of this intelligent selection process for few-shot examples, we compare the performance of GPT-4.1, which table 4 shows is the strongest model in this setting, against its performance with randomly selected few-shot examples in table 3. Although the model delivers strong performance when conditioned even on random exemplars, its performance is substantially improved through the intelligent exemplar retrieval process.
@@ Discussion
Interestingly, our experiments aggregated in fig. 3 reveal a notable divergence between models optimized for reasoning and for chat applications. Although reasoning models perform better at direct solving and zero-shot parsing into symbolic representations of rules and facts, the chat models exhibit surprising effectiveness at few-shot parsing of case facts. It is possible that reasoning models’ post-training focuses their efforts on emulating explicit reasoning steps, such as those Prolog would itself execute, rather than accurately mapping input texts into symbolic form. Whereas long chains of thought are useful for the complicated arithmetic required for directly calculating a tax obligation or for discerning which portions of the statutes should be included in a zero-shot parse, they degrade performance on the relatively simpler task of imitating exemplary conversions from natural language into formal logic. This could be a manifestation of the “lost in the middle” effect in long-context processing (Liu et al. 2024): when the model is producing its final parse, the case’s plain-text is in between a long block of the statutes’ Prolog and the long chain of thought which the model has just generated, thus distracting the model from the relatively straightforward task of converting from natural language into the neo-Davidsonian Prolog syntax which can comprehensively represent the SARA cases.
Figure 5: Cases with more words are more likely to be miscalculated. We note a moderate correlation between the case length and the failure rate of solution attempts.
In practice, this additionally means that with gold exemplars of case and statute translations, an individual’s tax case could be processed much more quickly than it could be without these manual annotations, because chat models take dramatically less time than reasoning models to produce answers. This could help build more interactive tax assistance tools which allow users to quickly iterate to amend and clarify their relevant tax information. This approach notably reduces the expected real-world cost per successful tax filing, highlighting the critical role that intelligent exemplar selection can play in guiding model outputs toward faithful and structured reasoning. The most effective configuration at minimizing the system’s break-even price is achieved when providing models with intelligently retrieved precedential examples, and deferring to a tax expert unless this system yields the same answer from two independently sampled solutions. The break-even price of \$49.48\\$49.48 for this system would save the average American taxpayer over $80\%$ of the amount spent on tax filing (Internal Revenue Service 2025). This strong performance could be a strong first step towards real-world pilot studies.
@@
We look closely at the cases which are most and least commonly solved correctly by the most powerful models (DeepSeek and OpenAI). We see in fig. 5 that cases which have more words in them are more likely to be mistaken across models, perhaps because they are more likely to contain intricate logical structures which are more difficult to translate. For instance, table 5 shows that both case 91 and case 83 were miscalculated by every approach. These two cases have an unusual predicate structure that does not appear in other cases, involving the person in question hiring overlapping groups of people for different periods of the year. Although either case appears in the other case’s few-shot examples when these are selected by the intelligent reranking system, the single example appears insufficient for any of the models we test to grasp the application of this unusual logical structure. Real-world deployment should consider how additional heuristics for case difficulty can further inform the decision to defer to a tax professional.
Least Frequently Failed Most Frequently Failed
Case Index Errors Case Index Errors
tax_case_89 0 tax_case_26 12
tax_case_11 1 tax_case_47 12
tax_case_76 1 tax_case_10 12
tax_case_45 1 tax_case_91 12
tax_case_56 2 tax_case_83 12
Table 5: Most and least frequently mistaken cases.
Finally, we note in table 2 a dramatic jump in performance with scale at parsing cases and statutes without gold translations. Although smaller reasoning models (those derived from Qwen2.5 32B and Llama 3.3 70B) perform solidly at direct solving, and the smaller chat models perform solidly at few-shot parsing, neither category of smaller model solves more than a handful of cases in the zero-shot parsing setting. In contrast, all larger models solve at least $10\%$ of these cases correctly, while for OpenAI o3 this is the most successful setting. We hope that further increases in model scale will further improve performance in this setting, since it combines the low human cost of direct solving with the easy auditing of systems that rely on the symbolic solver.

### v3

# Language Models and Logic Programs for Trustworthy Tax Reasoning

@@ Abstract
According to the United States Internal Revenue Service, “the average American spends $\mathdollar 270$ and 13 hours filing their taxes”. Even beyond the U.S., tax filing requires complex reasoning, combining application of overlapping rules with numerical calculations. Because errors can incur costly penalties, any automated system must deliver high accuracy and auditability, making modern large language models (LLMs) poorly suited for this task. We propose an approach that integrates LLMs with a symbolic solver to calculate tax obligations. We evaluate variants of this system on the challenging StAtutory Reasoning Assessment (SARA) dataset, and include a novel method for estimating the cost of deploying such a system based on real-world penalties for tax errors. We further show how combining up-front translation of plain-text rules into formal logic programs, combined with intelligently retrieved exemplars for formal case representations, can dramatically improve performance on this task and reduce costs to well below real-world averages. Our results demonstrate the effectiveness of applying semantic parsing methods to statutory reasoning, and show promising economic feasibility of neuro-symbolic architectures for increasing access to reliable tax assistance. Code is available at https://github.com/wjurayj/legal_logic_programs
@@ Introduction
Accuracy in tax filing is essential. Over-reported income or missed deduction opportunities lead to unnecessary over-payment, while under-reporting may result in penalties, interest, and potential legal consequences. In the United States, the costs of inaccuracies affect lower income communities more significantly, in part because these groups offer the Internal Revenue Service (IRS) high audit success rates (Dean 2022; Black et al. 2022; Elzayn et al. 2025). However, these audits deliver a modest return on investment compared to audits of wealthier taxpayers, so there is an opportunity to better align community and institutional interests through improved tax advice to lower income taxpayers (Boning et al. 2024).
@@
Our experiments are the first to demonstrate strong performance by applying semantic parsing methods to statutory reasoning. This includes two key findings. First, whereas frontier reasoning models outperform non-reasoning models at both directly solving and at parsing case and statute text into the symbolic solver, non-reasoning models consistently outperform their reasoning counterparts when given gold symbolic representations of statutes and of their application to similar cases. Second, we show that by adding additional refusal criteria through a symbolic solver and self-checking, the expected costs of deploying such a system in the real world could be brought down to less than 20% of the average cost for an American to file their taxes. Our results indicate the promise of neuro-symbolic architectures for expanding access to trustworthy and reliable tax expertise.
@@
Several programming languages have been designed to represent and facilitate logical reasoning. Prolog is a declarative programming language for representing and reasoning over knowledge, with roots in first order logic. A programmer defines rules using Horn clauses (Horn 1951) and facts by declaring which rules apply to entities, thus populating a knowledge base. Subsequently, this knowledge base can be queried by defining a ‘goal’, which launches computation in the form of a backward-chaining search attempting to prove that the goal holds a certain value by iteratively unifying sub-goals (Wielemaker et al. 2010; Weir et al. 2024). Prolog has been used since the early days of legal AI, where it has formed the backbone of legal expert systems because its declarative syntax keeps knowledge base entries for rules human-readable while powerfully representing the reasoning around which legal questions revolve (Sherman 1989). Efforts in countries like the United Kingdom (Sergot et al. 1986), Canada (Sherman 1987), and the United States (Kant et al. 2025) have leveraged this capacity to encode legal rules in executable formal logic. Related languages have also been used in the legal domain, such as Answer-set Programming (Gelfond and Lifschitz 1988; Morris 2020), Datalog (Ceri et al. 1989; Huang et al. 2021), and Catala (Merigoux et al. 2021a; Merigoux 2023). More broadly, hierarchical templates are a popular tool for evaluation of legal reasoning (Hou et al. 2024). We focus on the SARA dataset (Holzenberger et al. 2020), which encodes statutes and cases into Prolog logic programs to show how a symbolic expert system can perfectly solve a task which large language models struggle to complete.
@@
We evaluate our methods on the SARA dataset, which tests the ability of language models to do statutory reasoning about the United States Tax Code (Holzenberger et al. 2020). This dataset is included in the popular aggregate benchmark LegalBench (Guha et al. 2023), and was used in the GPT-4 product launch to highlight the model’s superior reasoning capacity (Blair-Stanek et al. 2024). The SARA dataset consists of 9 sections from the US federal tax code which have been moderately edited to make them self-contained and unambiguous. These manipulations allow the dataset to serve as a self-contained and solvable task for language models that lack live internet access and human-like abilities to process ambiguity (Jurayj et al. 2022; Stengel-Eskin et al. 2024). These statute sections are accompanied by 376 hand-crafted cases to test understanding of these statutes, each containing a question about a person’s tax obligation. Each statute and case has been manually translated into Prolog, which allows them to be trivially solved using the language’s powerful execution engine to resolve queries about cases. This Prolog is defined using neo-Davidsonian event semantics (Davidson 1966), categorizing each event as one of 61 possible predicates onto which various arguments are attached. 276 of these cases require binary responses about whether a section applies, and the remaining 100 require calculating a tax obligation; we focus on the 100 cases because of their increased difficulty and because a trivial baseline of always guessing a single answer delivers poor performance at predicting the numerical output.
Methodology
@@
Zero-Shot Parsing for a Symbolic Solver
To extend the direct solution approach, we augment the language model with a symbolic solver. Here, a model is given the plain text of the statutes as in the direct calculation case. It is instructed to generate a Prolog program which encodes the relevant rules and facts necessary to compute the person in question’s tax obligation, reframing the task from reasoning to autoformalization (Zhang et al. 2025a). The symbolic solver ingests a set of rules and facts in Prolog, and is invoked to execute a query, offloading to the solver the compositional reasoning that language models struggle to conduct faithfully (Khandelwal and Pavlick 2025). The execution of this Prolog program offers a straightforward mechanism for refusal: if the program fails to execute into the proper format or hangs beyond a pre-allocated time limit (10 seconds), the system is considered to have refused to answer.
@@
Table 2: Results of different methods without gold statutes. Models in the same family have the same base model (or seem most likely to, in the case of closed-weights models). Note that “break-even price” measures only the costs of failures and abstentions, and does not include inference costs. For each model, the approach that delivered the lowest break-even price is shown in bold. The top two rows show the break-even price of trivial systems, which always defer to an expert or which always tell a person not to pay any taxes. Errors represent a symmetricized 90% confidence interval. The lowest break-even price method for each model family is in bold.
Few-Shot Parsing using Gold Statutes
We additionally consider what advantage could be given by offering language model parsers access to gold symbolic representations of the statutes, alongside demonstrations of cases applying the rules from these statutes, building on previous work showing how demonstrations can aid in semantic parsing (Shin and Van Durme 2022; Spiegel et al. 2024). Notably, the gold parsed cases all reference the same manually-translated Prolog representation of the statutes. As such, these manual translations formalize certain types of facts in one particular way where several may be viable, given alternative representations of the rules which are plausible but not implemented in practice. Thus, when using already parsed cases as few-shot examples, the execution engine must have access to this particular formalization of the rules as well. This reduces the per-case reasoning task to tax-relevant event extraction (Holzenberger and Van Durme 2023; Gantt et al. 2024; Walden et al. 2025)
To identify the most salient exemplar cases, we apply information retrieval systems which can follow instructions (Weller et al. 2025a), directing a retrieval system to rank cases based on how similar the logical structure of the case’s text is to the case at hand. For each case, we instruct a lightweight reasoning model (OpenAI o4-mini) to rank the other 99 cases, following recent work showing the effectiveness of test-time scaling for reranking documents (Weller et al. 2025b; Yang et al. 2025). As few-shot examples, we provide the 5 most relevant cases and their gold Prolog translations in-context for the language model to condition its parse on. These cases can be analogized to precedents, because our tax calculation agent uses them to understand how terms from the statutes are applied in practice. We note that this use of the term ‘precedent’ is informal and does not refer to the common-law practice of binding precedent. Rather, these retrieved precedent cases are more analogous to ‘persuasive’ precedent, which might help a court understand how terms have been applied in the past but is not itself new law which future courts must apply (Kozel 2014; Savelka 2020; Šavelka and Ashley 2022).
@@
We run experiments across four model families of different sizes, three of which are open-weight models. The bases for these models are: Qwen2.5 32B (Qwen et al. 2025), Llama 3.3 70B (Grattafiori et al. 2024), DeepSeek-V3 671B (DeepSeek-AI et al. 2025b), and OpenAI’s GPT-4.1 (OpenAI 2025a, b), each of which has an instruction-tuned version designed for common chat applications, and a reasoning version optimized to expend additional inference-time compute to solve harder problems. The full list of models is included in table 1. We run auxiliary experiments using GPT-5, but we do not conduct the same chat vs. reasoning comparison because this product appears to be an integrated system containing several models, and therefore is not as analogous to these other comparisons (Zhang et al. 2025b). The reasoning model for three of these pairings stem from the DeepSeek R1 project (DeepSeek-AI et al. 2025a), where strong base models were fine-tuned to generate long chains-of-thought that help them solve harder quantitative reasoning problems. Although there is no formal documentation stating that OpenAI’s GPT-4.1 and o3 models are derived from the same base model, the proximity of the two model’s launch dates and their identical per-token pricing schemes suggest that these two models have a similar relationship to the other model pairs we explore.
@@
We further ask how effectively these methods can serve to improve each other’s selectivity, by using comparisons between different solution methods to expend additional compute to help determine whether an answer should be trusted (Wang et al. 2023; Stengel-Eskin and Van Durme 2023). In these settings, an answer is only accepted if it is reached via two independent reasoning processes: two chains-of-thought and answers (either directly calculated obligations or Prolog programs) are sampled from the same model. When self-checking using the same method (for instance, “Parsed + Parsed”), these answers are conditioned on the same prompt and context. In the parsing-based approaches, this can be considered a more stringent version of the existing refusal system which rejects questions for which the parsed solution does not execute, by additionally deferring to a tax professional where a combination of attempts do not reach a consensus answer. We test the effectiveness of each combination of reasoning processes for each model to show where additional selectivity can further improve performance.
@@
We posit that the deployment of an automated tax advisor system should require some accountability for the organization deploying it. To provide a realistic estimate of the costs of employing the methods we outline above draw from the US Internal Revenue Code (IRC) $\S 6662$, which imposes a penalty of $20\%$ of the amount underpaid for “substantial understatement of income tax”. The cases in SARA are all personal or household tax cases, so a tax filing is considered a “substantial understatement” if they understate the true obligation by more than the maximum of $10\%$ of the actual amount owed or $\mathdollar 5,000$. We finally impose a penalty of $\mathdollar 270$ for refusing to answer, to model the cost of the personal or professional time required to complete one’s taxes (Internal Revenue Service 2025).
@@
$\text{cost}=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}-(\Delta y_{i}),&\Delta y_{i}<0,\\[6.0pt] 0.2\,(\Delta y_{i}),&\Delta y_{i}>\max\!\bigl(\mathdollar 5,000,\;0.1\cdot y_{i}\bigr),\\[6.0pt] $\textdollar 270$,&\text{if refused},\\[6.0pt] $\textdollar 0$,&\text{otherwise}.\par\end{cases}$
@@ Discussion
Interestingly, our experiments aggregated in fig. 3 reveal a notable divergence between models optimized for reasoning and for chat applications. Although reasoning models perform better at direct solving and zero-shot parsing into symbolic representations of rules and facts, the chat models exhibit surprising effectiveness at few-shot parsing of case facts. It is possible that reasoning models’ post-training focuses their efforts on emulating explicit reasoning steps, such as those Prolog would itself execute, rather than accurately mapping input texts into symbolic form. Whereas long chains of thought are useful for the complicated arithmetic required for directly calculating a tax obligation or for discerning which portions of the statutes should be included in a zero-shot parse, they degrade performance on the relatively simpler task of imitating exemplary conversions from natural language into formal logic. This aligns with concurrent work highlighting cases in which long chains of thought can distract models from tasks where shorter chains would deliver stronger performance (Gema et al. 2025).
Figure 5: Standalone parsing success rate improves with model size. While smaller reasoning models fail to solve nearly every problem in this setting, larger reasoning models show improvement with model size.
In practice, this additionally means that with gold exemplars of case and statute translations, an individual’s tax case could be processed much more quickly than it could be without these manual annotations, because chat models take dramatically less time than reasoning models to produce answers. This could help build more interactive tax assistance tools which allow users to quickly iterate to amend and clarify their relevant tax information. This approach notably reduces the expected real-world cost per successful tax filing, highlighting the critical role that intelligent exemplar selection can play in guiding model outputs toward faithful and structured reasoning. The most effective configuration at minimizing the system’s break-even price is achieved when providing models with intelligently retrieved precedential examples, and deferring to a tax expert unless this system yields the same answer from two independently sampled solutions. The break-even price of $\mathdollar 15.78$ for this system would save the average American taxpayer over $90\%$ of the amount spent on tax filing (Internal Revenue Service 2025). This strong performance could be a strong first step towards real-world pilot studies, or on developing open-source model alternatives to GPT-5 that can adaptively allocate computational resources based on problem difficulty (Jurayj et al. 2025; Wang et al. 2025, 2026).
@@
Furthermore, we believe the relative strength of o3 at zero-shot parsing (which includes parsing a subset of the statutory rules) suggests promise in semi-automated offline encoding of these rules; as models improve further at this type of complex semantic parsing, entire legal systems could be distilled into programmatic logic. We note in fig. 5 a dramatic jump in performance with scale at parsing cases and statutes without gold translations. Although smaller reasoning models (those derived from Qwen2.5 32B and Llama 3.3 70B) perform solidly at direct solving, and the smaller chat models perform solidly at few-shot parsing, neither category of smaller model solves more than a handful of cases in the zero-shot parsing setting. In contrast, all larger models solve at least $10\%$ of these cases correctly, while for OpenAI o3 this is the most successful setting. We hope that further increases in model scale will further improve performance in this setting, since it combines the low human cost of direct solving with the easy auditing of systems that rely on the symbolic solver. Additionally, this capability could also align with improved performance on formalizing statutes alone, allowing tighter feedback for policy makers on the effects of proposed amendments on government revenue (Blair-Stanek et al. 2026).